Karen runs a print shop that makes posters for large companies. It is a very competitive business. The market price is currently $1 per poster. She has fixed costs of $250. Her variable costs are $1,800 for the first thousand posters, $1,500 for the second thousand, and then $900 for each additional thousand posters.



Instructions: Round your answers to 3 decimal places.



a. What is her AFC per poster (not per thousand!) if she prints 1,000 posters? $
.



What if she prints 2,000 posters? $
.



What if she prints 10,000 posters? $
.



b. What is her ATC per poster if she prints 1,000? $
.



What if she prints 2,000? $
.



What if she prints 10,000? $
.



c. If the market price fell to 85 cents per poster, would there be any output level at which Karen would not shut down production immediately? .
1
Expert's answer
2020-05-08T14:15:42-0400

p=1, FC=250


a)

"AFC= \frac {FC}{Q}=\frac {250}{1000}=0.25"

"AFC=\frac{FC}{Q}=\frac {250}{2000}=0.125"


"AFC=\frac{FC}{Q}=\frac {250}{10000}=0.025"

b)


"ATC=\frac {TC}{Q}=\frac {FC+VC}{Q}=\frac {250+1800}{1000}=2.05"


"ATC=\frac{250+1800+1500}{2000}=1.775"


"ATC=\frac {250+1800+1500+900 \times 8}{10000}=1.075"

c)


"\frac {250+1800+1500+900 \times n}{2000+n}=0.85"


"n<0"

This level of production will not be